OpenStudy (anonymous):

introduction to differential equations, calc one. introduction to differential equations, calc one. @Mathematics

OpenStudy (anonymous):

\[dz/dt = t^2z^2\]

OpenStudy (amistre64):

seperation of variables

OpenStudy (anonymous):

I multiplied both sides by Dtz^3, but now do I need to apply the product rule, and if so, how do I do that backwards?

OpenStudy (amistre64):

the z^3 is a little odd

OpenStudy (anonymous):

z^2, sorry.

OpenStudy (amistre64):

multiply boths sides by dt dt(dz/dt=z^2 t^2) dz = z^2 t^2 dt now divide both sides by z^2 (dz = z^2 t^2 dt)/z^2 dz/z^2 = t^2 dt , and integrate

OpenStudy (amistre64):

\[\int \frac{1}{z^2}dz=\int t^2\ dt\] \[-\frac{1}{z}+c_1=\frac{1}{3} t^3+c_2\] \[-\frac{1}{z}=\frac{1}{3} t^3+c_2-c_1\] \[-\frac{1}{z}=\frac{1}{3} t^3+C\]

OpenStudy (amistre64):

might have to do some minor mathing to get z in the proper condition yet, but thats the brunt of it

OpenStudy (anonymous):

my struggle is intergrating the z, I follow everything you have done with the t side.

OpenStudy (amistre64):

the z part then, might be better to see it as z^(-2) instead

OpenStudy (amistre64):

\[\int \frac{1}{z^2}dz=\int z^{-2}dz\]

OpenStudy (amistre64):

then the power rule for integration is the same as the t side .... add one and divide by it

OpenStudy (amistre64):

\[\int z^{-2}dx=\frac{z^{-2+1}}{-2+1}=-z^{-1}=-\frac{1}{z}\]

OpenStudy (anonymous):

roger, that's a big help, but let me ask you, why does c2-c1 turn into a singular C?

OpenStudy (amistre64):

Its just a convenient way to absorb the arbitrary constants into one mass of arbitrary constant-ness...

OpenStudy (amistre64):

when i first learned these i was confused as to why they didnt put a +c on the left side; i eventually found out why

OpenStudy (anonymous):

its rough, i thought dx/dy was a symbol until last week.

OpenStudy (amistre64):

yeah, they let you think that as well.

OpenStudy (amistre64):

its an operator

OpenStudy (amistre64):

like + is the operator for addition; / is the operator for divide, its just an operator

OpenStudy (amistre64):

at this level of math you start to see that math isnt as stringent and concrete as you were led to believe

OpenStudy (anonymous):

blue pill please.

OpenStudy (amistre64):

blue pill?